OpenStudy (anonymous):

Subtraction of Rational Expressions x/4 - y/3

OpenStudy (anonymous):

\[\frac{ x }{ 4 }-\frac{ y }{ 3 }\]

OpenStudy (isaiah.feynman):

\[\frac{ a }{ b }-\frac{ c }{ d } = \frac{ ad-bc }{ bd }\]

OpenStudy (anonymous):

But the book says a/c - b/c = a-b/c , so im confused...

OpenStudy (isaiah.feynman):

Oh in the book's case the denominators are the SAME.

OpenStudy (isaiah.feynman):